# Inference-time Scaling of Diffusion Models through Classical Search

### v1

@@ E.1 Long Horizon Maze Planning
For all our maze experiments we use the OGBench PointMaze [55]. PointMaze Giant was directly adapted from their datasets, and we create PointMaze Ultra using the same protocol of Figure 5 in [49] [^1] https://github.com/mpetersen94/gcs/blob/main/models/maze.py, but with a smaller size of 20x20 cells. Dataset for Ultra is collected following the protocal in OGBench [55]. A larger layout and example trajectory of PointMaze Ultra is shown in Fig. 7. We evaluate the model on the default task 1 of OGBench [55], which is navigating from bottom left to top right. Empirically we discover that the diffusion model can perform well on short-horizon tasks without extra inference compute, but struggles heavily in the long horizon tasks.
@@
We train the model following diffuser [31], where we use a temporal U-Net to denoise the trajectory. Since our objective start and goal is more distant than trajectories in dataset, we sample at longer horizons than training, which is enabled by the temporal U-Net architecture. Specifically, for PointMaze Giant, we train at horizon 400 and sample at 600; and for PointMaze Ultra, we train at horizon 2000 and sample at 2800. For the PointMaze Ultra model we extend the depth of the U-Net for 1 more layer, and keep all other parameters the same. We train the model for 1.2M steps using the same configuration as [31].
@@
For local search, we search the hyper-parameters $\bar{\rho}$ and $\bar{\mu}$ in Sec. C.5 with $\bar{\gamma}=0$, and we set the local search steps (recurrence steps) to $2$ for Giant and $6$ for Ultra. We set $N_{\text{iter}}=1$ as we observe they do not contribute to the performance gains, while the number of Langevin MCMC steps has the most impact on performance. For global search with BFS we evaluate at steps $\left\{12,8,4\right\}$, and for DFS we evaluate at $\left\{12,11,\cdots,1\right\}$ with backtracking depth $\Delta_{T}=4$. We also observe that increasing backtracking depth to $12$ and evaluate at smaller time steps $\left\{4,3,2,1\right\}$ helps to scale up the performance with more compute. We use grid search to determine temperature and threshold, with a step size of $[1,5,10]$ for each magnitude.
E.2 Offline RL
@@
We follow the setup in [44], and we directly use their pretrained diffusion model and Q-function, omitting the time-dependent energy function. The diffusion model was trained to generate action $\bm{a}$ given state $\bm{s}$, and we sample with 15 steps of DDIM.
For hyper-parameter search, we disable the implicit dynamics and set $\bar{\gamma}=0$, and use the ‘increase’ schedule for $\bm{\rho}$ and $\bm{\mu}$. For strength parameters $\bar{\rho}$ and $\bar{\mu}$, we first search for the right magnitude. Then, we also follow [44] and search with step size [1,2,3,5,8,10] within the magnitude. Same as [44], we use 5 different seeds with 10 samples per seed for each task. To avoid over fitting, we use different seeds for parameter search and evaluation. We report the hyper-parameters and the performance within the parameter-searching dataset and evaluation dataset.
For global search, we use 4 particles for Medium-Expert and Medium datasets, and 2 particles for Medium-Replay datasets. Since the number of particles are small, we do not carry out BFS or DFS methods and simply use Best-of-N. We point out that the number of particles we use are much smaller than the 50 particles in [80] and the 32 particles in [8], highlighting the effectiveness of local search.
@@
We compare our method to a variety of baselines, including traditional state-of-the-art methods like CQL [35], BCQ [20] and IQL [34], diffusion-based policies such as diffuser [31], decision-diffuser (DD) [1], Diffusion-QL (D-QL) [80], SfBS [8] and QGPO [44]. We directly take the numbers from [44].
Among the baseline diffusion-based methods, both Diffuser [31] and QGPO [44] requires training a noise-dependent guidance function, and D-QL [80] requires updating the diffusion model during training using the Q-function iteratively, which needs to back-propagate through the diffusion sampling chain, introducing high computation and memory overheads. DD [1] uses classifier-free guidance [27] to generate high-return trajectories, which requires training a return-conditional model on labeled datasets, which can be expensive to obtain in robotics where only demonstration data is available [43].
@@
E.3 Image Generation
E.3.1 Text-to-Image Compositional Generation
@@
We use the SSD-1B model [^2] https://huggingface.co/segmind/SSD-1B which is distilled from SDXL, and we use the default sampling configuration with 50 steps of DDIM sampler. We use the oracle verifier of each task following CompBench [29].
Global search with DFS
For DFS we evaluate at time steps $\left\{25,35,45\right\}$ and set the backtrack depth $\Delta_{T}=25$. We set the threshold for our experiments to $0.5$, $0.7$ and $0.9$, with the ‘constant’ schedule in Eq. 14. We typically use a higher threshold for high computation budgets, representing using more compute for higher quality standard.
We provide example visualizations in Fig. 8.
(a) 1 Particle
(b) 2 Particles
(c) 4 Particles
(d) 8 Particles
Figure 8: Example of inference-time scaling with Numeracy dataset Above shows the results of Best-of-N with different particles, with prompt: “four dogs played with two toys”
E.3.2 ImageNet Conditional Generation
Setup
We use the 256x256 DDPM ImageNet model from [12], and for local search we use the 224px ViT classifier [^3] https://huggingface.co/google/vit-base-patch16-224 and for global search we use the 384px version [^4] https://huggingface.co/google/vit-base-patch16-384 when using double-verifier. We evaluate on class label 222 and 444, with 256 samples per label. We sample with 100 DDPM steps. When evaluating the class accuracy we use the DeiT model [^5] https://huggingface.co/facebook/deit-small-patch16-224, and we evaluate the FID againt the images of the corresponding class in ImageNet.
Global search with BFS
Due to computation constraints we set the local search steps to $1$ in all experiments. For BFS we set the evaluation steps as $\left\{75,50,25\right\}$, and use $\bar{\tau}=1.0$ for label 222 and $\bar{\tau}=0.5$ for label 444 with the ‘increase’ schedule.
We provide example visualizations in Fig. 9.
(a) 2 Particles
(b) 4 Particles
(c) 8 Particles
(d) 12 Particles
Figure 9: Visualizations of ImageNet Conditional Generation Above shows the uncurated samples for class label 222 with BFS with different number of particles

### v2

@@ E.3 Long Horizon Maze Planning
For all our maze experiments we use the OGBench PointMaze environment (Park et al., 2024). We created our maze layout using the same protocol of Figure 5 in Marcucci et al. (2023) [^2] https://github.com/mpetersen94/gcs/blob/main/models/maze.py, but with a smaller size of 20x20 cells. Dataset is collected following the protocal in OGBench (Park et al., 2024). We evaluate the model on the default task 1 of OGBench (Park et al., 2024), which is navigating from bottom left to top right. Empirically we discover that the diffusion model can perform well on short-horizon tasks without extra inference compute, but struggles heavily in the long horizon tasks.
@@
We train the model following diffuser (Janner et al., 2022), where we use a temporal U-Net to denoise the trajectory
Since our objective start and goal is more distant than trajectories in dataset, we sample at longer horizons than training, which is enabled by the temporal U-Net architecture. We train the model for 1.2M steps using the same configuration as Janner et al. (2022). Model details are below in Table. 7.
Table 7: Hyper-parameters for maze model
hidden dimension multipliers [1,4,8,16]
training horizon 2000
sampling horizon 2800
training steps 1.2M
batch size 64
learning rate 1e-4
@@
For local search, we search the hyper-parameters $\bar{\rho}$ and $\bar{\mu}$ in Sec. C.5 with $\bar{\gamma}=0$. For global search with BFS we evaluate at steps $\left\{12,8,4\right\}$, and for DFS we evaluate at $\left\{12,11,\cdots,1\right\}$ with backtracking depth $\Delta_{T}=4$. We also observe that increasing backtracking depth to $12$ and evaluate at smaller time steps $\left\{4,3,2,1\right\}$ helps to scale up the performance with more compute. The hyperparameter search results are below:
N $\tau=0.2$ $\tau=0.005$ $\tau=0.1$
2 $27.5\pm 4.3$ $32.5\pm 1.1$ $31.2\pm 4.2$
4 $42.5\pm 5.2$ $48.1\pm 1.1$ $45.5\pm 2.3$
8 $67.6\pm 1.1$ $71.2\pm 2.2$ $70.1\pm 1.1$
Table 8: Hyperparameter search for temperature $\tau$ in PointMaze BFS
Figure 8: Hyperparameter search for threshold $\delta$ in PointMaze DFS
E.4 Offline RL
Background
Diffusion policy (Chi et al., 2023) is widely used for action generation in robot foundation models (Black et al., 2024; Liu et al., 2024). At inference time, policies can be guided by human trajectory constraints (Wang et al., 2022) or LLM-based value functions (Nakamoto et al., 2024). Exact sampling requires training a noise-dependent energy function (Lu et al., 2023), but this can degrade pretrained knowledge and demands additional data—often impractical in data-scarce robotic settings. In contrast, inference-scaling provides a more flexible approach, allowing seamless composition of pretrained diffusion policies with Q-functions without retraining.
@@ E.4 Offline RL
We follow the setup in Lu et al. (2023), and we directly use their pretrained diffusion model and Q-function, omitting the time-dependent energy function. The diffusion model was trained to generate action $\bm{a}$ given state $\bm{s}$, and we sample with 15 steps of DDIM.
For hyper-parameter search, we disable the implicit dynamics and set $\bar{\gamma}=0$, and use the ‘increase’ schedule for $\bm{\rho}$ and $\bm{\mu}$. For strength parameters $\bar{\rho}$ and $\bar{\mu}$, we first search for the right magnitude. Then, we also follow Lu et al. (2023) and search with step size [1,2,3,5,8,10] within the magnitude. Same as Lu et al. (2023), we use 5 different seeds with 10 samples per seed for each task. To avoid over fitting, we use different seeds for parameter search and evaluation. We report the hyper-parameters and the performance within the parameter-searching dataset and evaluation dataset.
For global search, we use 4 particles for Medium-Expert and Medium datasets, and 2 particles for Medium-Replay datasets. Since the number of particles are small, we do not carry out BFS or DFS methods and simply use Best-of-N. We point out that the number of particles we use are much smaller than the 50 particles in Wang et al. (2022) and the 32 particles in Chen et al. (2023), highlighting the effectiveness of local search.
@@
We compare our method to a variety of baselines, including traditional state-of-the-art methods IQL (Kostrikov et al., 2021) and diffusion-based policies such as diffuser (Janner et al., 2022), decision-diffuser (DD) (Ajay et al., 2022), Diffusion-QL (D-QL) (Wang et al., 2022), SfBC (Chen et al., 2023) and QGPO (Lu et al., 2023). We directly take the numbers from Lu et al. (2023).
Among the baseline diffusion-based methods, both Diffuser (Janner et al., 2022) and QGPO (Lu et al., 2023) requires training a noise-dependent guidance function, and D-QL (Wang et al., 2022) requires updating the diffusion model during training using the Q-function iteratively, which needs to back-propagate through the diffusion sampling chain, introducing high computation and memory overheads. DD (Ajay et al., 2022) uses classifier-free guidance (Ho & Salimans, 2022) to generate high-return trajectories that requires training a return-conditional model on labeled datasets, which can be expensive to obtain in robotics where only demonstration data is available (Liu et al., 2024).
For our reproduced baselines, TFG (Ye et al., 2024) is allowed up to 8 recurrence steps and DAS (Kim et al., 2025) up to 16 particles, resulting in a hyperparameter space and computational cost approximately twice that of our method. We sweep across all configurations for the baseline methods and report the best performance. For fair comparison we evaluate our method on different seeds used for hyperparameter search, with the results shown in Table. 9.
@@
E.4.1 Offline Policy Distillation
@@
We adopt the Medium dataset from D4RL (Fu et al., 2020) and the corresponding pretrained models from Lu et al. (2023) as in the previous section. For the dataset, we replace the action of each state with the sample generated with test-time search (TTS), using the hyper-parameters in Table. 9. We then finetune the models on the modified dataset with early-stopping. For evaluation of finetuned models, we use 5 different seeds different from the seed used during training, with 10 samples per seed. During evaluation, we sample from the finetuned diffusion model directly, without the presence of any verifier. Detailed training hyper-parameters are below in Table. 10.
epochs 5
learning rate 1e-4
batch size 16384
eval every epoch 1
sampling steps 15
Table 10: Hyper-parameters for policy distillation
For the Baseline DPPO (Ren et al., 2024), we directly adopt the numbers reported in their paper, which uses the same Medium dataset and models as ours.
The training curves for policy finetuning are shown in Fig. 9. In both Hopper and Walker2d environments, performance converges rapidly, within a single epoch over the dataset. Notably, our method is significantly faster than online finetuning approaches, as it does not require online data collection. Training one epoch on the full Medium dataset takes only about 4 minutes on a single Nvidia RTX 4090 GPU, including the time for data generation via inference-scaling sampling, whereas DPPO requires several hours to reach convergence.
Figure 9: Training curves for policy distillation
E.5 Mitigating reward hacking with double verifier
Experimental setup
We evaluate the proposed double-verifier on the challenging conditional ImageNet generation task, generating target-class samples from an unconditional model guided by a pretrained classifier. Specifically, we use two independent classifiers as verifiers [^3] https://huggingface.co/google/vit-base-patch16-224444https://huggingface.co/google/vit-base-patch16-384 for global and local search. We report the Fréchet Inception Distance (FID) computed on 256 generated samples against the corresponding ImageNet class, and measure class accuracy using a separate classifier [^5] https://huggingface.co/facebook/deit-small-patch16-224. Since we only apply the global verifier sparsely, double-verfier introduces negligible computational costs. For local search we search the hyper-parameters with step size [1,2,4,8,10], and for global search with BFS, we search the temperature in [1.0,2.0,4.0]. For MSP score, we adopt a different imagenet classifier  [^6] https://huggingface.co/facebook/deit-base-patch16-224.
For the MSP score (Hendrycks & Gimpel, 2016), it is defined as:
$\text{MSP}({\bm{x}})=\max_{c}p(c|{\bm{x}})$
Thus, higher MSP score indicates higher confidence that the image belongs to one of the classes from the dataset, which is less OOD.
Visualizations
Here we provide some visualizations of adversarial reward hacking and the effectiveness of our double-verifier approach. As shown in Fig. 10, we can see the adversarial patches that exploits the weakness of a single verifier, making the verifier to predict the images as belonging to the wrong class.
(a) Adversarial example with 0.96 class validity predicted by the local search verifier
(b) Adversarial example with 0.95 class validity predicted by the local search verifier
(c) Adversarial example with 0.98 class validity predicted by the local search verifier
Figure 10: Examples of adversarial samples generated by gradient over-optimization Here we provide the adversarial samples generated by gradient guidance on class 222 (kuvasz). Although the samples have high probability of belonging to the target class as predicted by the verifier, they are caused by adversarial reward hacking.